OpenStudy (bacon):

What is the percent of discount on an item that has a sale price of $179.95 and regular price of $213.95? Round to the nearest whole number percent.

OpenStudy (anonymous):

Similar... Imagine starting price of $20, and sale price of $15. Discount is $5, but what is % ?

OpenStudy (anonymous):

i like your name, bacon

OpenStudy (anonymous):

you calculate the discount % as a comparison of the actual discount in $ versus the starting price.

OpenStudy (anonymous):

Everything's better with... @Bacon

OpenStudy (bacon):

@HorribleAtMath thanks, il like yours too! :D

OpenStudy (bacon):

hmmmm?

OpenStudy (anonymous):

Did you follow my comment above about calculating the discounts?

OpenStudy (bacon):

the first one and the second one yeah

OpenStudy (anonymous):

Sale: $179.95 regular $213.95 Discount amount = 213.95 - 179.95 = 34.00 Discount % = 34/213.95

OpenStudy (bacon):

16%

OpenStudy (bacon):

sooo

OpenStudy (bacon):

for...... What is the percent of discount on an item that has a sale price of $675.05 and regular price of $750.50? Round to the nearest whole number percent.

OpenStudy (bacon):

it would be 750.50-675.05=75.45

OpenStudy (bacon):

75.45/750.50

OpenStudy (bacon):

10%

OpenStudy (anonymous):

Looks right to me (if you rounded that right)

OpenStudy (bacon):

ok thanks for all the help

OpenStudy (bacon):

if you want another meddle for that tell me

OpenStudy (anonymous):

heh we all love the medals :) did you already do it on this one? If so, thanks and don't worry about it :) I enjoy helping :)

OpenStudy (bacon):

i can open a new one

OpenStudy (anonymous):

do you have another question?

OpenStudy (anonymous):

oinkity oink oink oink

OpenStudy (bacon):

LOL

OpenStudy (bacon):

oinkety oinkety oinkoink